# Causal Structure Learning in Hawkes Processes with Complex Latent Confounder Networks

### v1

@@ 3.2 Structure Discovery Through Rank Constraints
Consider a PO-MHP with excitation function $\phi_{ij}(s)=a_{ij}w(s)$ and rank faithfulness. The system $\mathcal{N}_{\mathcal{G}}\coloneq\mathcal{O}_{\mathcal{G}}\cup\mathcal{L}_{\mathcal{G}}$ consists of observed subprocesses $\mathcal{O}_{\mathcal{G}}\coloneq\{O_{i}\}_{i=1}^{p}$, and inferred latent confounder processes $\mathcal{L}_{\mathcal{G}}$ whose parent cause sets are yet to be identified. Let $\mathbf{O}_{v}\coloneq\{O_{i}^{(j)}\}^{j\in\{n-m,\dots,n\}}_{i\in\{1,\dots,p\}}$ denote the corresponding observed variable set. For a subprocess $N_{1}\in\mathcal{N}_{\mathcal{G}}$ and a candidate parent cause set $\mathcal{P}^{\prime}_{\mathcal{G}}\subseteq\mathcal{N}_{\mathcal{G}}$, when either $N_{1}$ is latent, or $\mathcal{P}^{\prime}_{\mathcal{G}}$ contains latent subprocesses, or both, the following condition holds: $\mathcal{P}^{\prime}_{\mathcal{G}}$ is the minimal set such that $\mathrm{rank}\left(\Sigma_{\mathbf{A}_{v},\mathbf{B}_{v}}\right)=|\mathbf{A}_{v}|-1$, where $\mathbf{A}_{v}\coloneq\{\mathcal{D}e(N_{1})^{(j)},\mathcal{D}e(L_{i})^{(j)}\}^{j\in\{n-m,\dots,n\}}_{L_{i}\in\mathcal{P}^{\prime}_{\mathcal{G}}}\cup\{O_{i}^{(j)}\}^{j\in\{n-m,\dots,n-1\}}_{O_{i}\in\mathcal{P}^{\prime}_{\mathcal{G}}}\cup\{O_{i}^{(j)}\}^{j\in\{n-m,\dots,n\}}_{O_{i}\in\mathcal{S}ib(\mathcal{D}e(N_{1}))\,\cup\,\{\mathcal{S}ib(\mathcal{D}e(L_{i}))\}_{L_{i}\in\mathcal{P}^{\prime}_{\mathcal{G}}}}$ and $\mathbf{B}_{v}=\mathbf{O}_{v}\backslash\left(\mathcal{D}e(N_{1})^{(n)}\cup\{\mathcal{D}e(L_{i})^{(n)}\}_{L_{i}\in\mathcal{P}^{\prime}_{\mathcal{G}}}\right)$ , if and only if $\mathcal{P}^{\prime}_{\mathcal{G}}$ is a subset of the parent cause set of $N_{1}$ such that: conditioning on $\mathcal{S}_{\mathcal{G}}\coloneq\mathcal{P}^{\prime}_{\mathcal{G}}\cup\mathcal{D}e(N_{1})\cup\{\mathcal{D}e(L_{i})\}_{L_{i}\in\mathcal{P}^{\prime}_{\mathcal{G}}}\cup\mathcal{S}ib(\mathcal{D}e(N_{1}))\cup\{\mathcal{S}ib(\mathcal{D}e(L_{i}))\}_{L_{i}\in\mathcal{P}^{\prime}_{\mathcal{G}}}$ renders $N_{i}$ locally independent of $\mathcal{N}_{\mathcal{G}}\backslash\mathcal{S}_{\mathcal{G}}$; for each $L_{i}\in\mathcal{P}^{\prime}_{\mathcal{G}}$, the latent confounder $L_{i}$ with observed effects $\{\mathcal{D}e(N_{1}),\mathcal{D}e(L_{i})\}$ satisfies Definition˜3.4; and, all possible observed surrogates of $N_{i}$ in $\mathcal{O}_{\mathcal{G}}$ have been identified so as to be added into the observed sibling set.
@@
With Theorem˜3.7 (and Proposition˜3.3), we can identify arbitrary causal relations among both observed and inferred latent subprocesses. This naturally raises a final question: How can we further infer new latent subprocesses that are causally related to inferred latent subprocesses, as in Fig. 3(d)? As shown in the following theorem, the observed surrogate of a latent subprocess can still be leveraged for such inference.
@@
Theorem˜3.7 and Theorem˜3.8 are extensions of Proposition˜3.3 and Proposition˜3.5, respectively. These extend the framework by replacing latent subprocesses with their observed surrogates when evaluating the rank of the relevant sub-covariance matrices. Equipped with these four key theorems, we are now ready to present the discovery algorithm in the next section.
@@ 4 Rank-based Discovery Algorithm
In this section, we present a two-phase iterative algorithm that leverages the identification theorems to iteratively identify causal relationships among discovered subprocesses and discover new latent subprocesses. Let $\mathcal{A}_{\mathcal{G}}$ denote the active process set, consisting of subprocesses whose parent causes are yet to be identified. Initially, $\mathcal{A}_{\mathcal{G}}$ is set to the observed subprocess set $\mathcal{O}_{\mathcal{G}}$ and is progressively updated throughout the procedure. Additionally, due to the existence of cycles in the summary causal graph, observed subprocesses previously identified as effects may still serve as causes for other subprocesses in $\mathcal{A}_{\mathcal{G}}$, and thus remain under investigation. The overall procedure is outlined in Algorithm˜1.
@@
1: Initialize partial causal graph $\mathcal{G}\coloneq\emptyset$, active process set $\mathcal{A}_{\mathcal{G}}\coloneq\mathcal{O}_{\mathcal{G}}$.
2: Exclude all observed subprocesses that are unconditionally locally independent of the others, such as single-node subprocesses, by testing unconditional independence relations among the corresponding variables.
3:repeat
4:  $(\mathcal{G},\mathcal{A}_{\mathcal{G}})\leftarrow$ Identifying Causal Relations $(\mathcal{G},\mathcal{A}_{\mathcal{G}},\mathcal{O}_{\mathcal{G}})$.  //phase I
5:  $(\mathcal{G},\mathcal{A}_{\mathcal{G}})\leftarrow$ Discovering New Latent Subprocesses $(\mathcal{G},\mathcal{A}_{\mathcal{G}},\mathcal{O}_{\mathcal{G}})$.  //phase II
6:until $\mathcal{A}_{\mathcal{G}}$ is empty or no updates occur.
7:return: $\mathcal{G}$
@@
Each iteration begins with Phase I, which aims to identify the causal structure for under-investigated subprocesses (both latent and observed) in $\mathcal{A}_{\mathcal{G}}$. In this phase, we systematically iterate over each subprocess in $\mathcal{A}_{\mathcal{G}}$ and attempt to identify its parent causes using the current $\mathcal{A}_{\mathcal{G}}\cup\mathcal{O}_{\mathcal{G}}$. If a subprocess’s parent cause set is fully contained within this set, it can be identified using Propositions˜3.3 and 3.7. Once its parent cause set is identified, the subprocess is removed from $\mathcal{A}_{\mathcal{G}}$. This phase continues until no further updates occur. Details of this phase are provided in Algorithm˜2 in Section˜O.1.
@@
When no more subprocesses in $\mathcal{A}_{\mathcal{G}}$ can be resolved using Phase I, we enter Phase II. This phase seeks to discover new latent confounder subprocesses by exhaustively checking all pairs in $\mathcal{A}_{\mathcal{G}}$ using Propositions˜3.5 and 3.8. Identified latent confounders are merged if they overlap in subprocesses, implying they share the same latent parent cause. $\mathcal{A}_{\mathcal{G}}$ is then updated to add new latent subprocesses and remove their effects, and the algorithm returns to Phase I in the next iteration. The procedure continues until $\mathcal{A}_{\mathcal{G}}$ is empty or remains unchanged. Detailed steps are provided in Algorithm˜3 in Section˜O.2.
@@
Consider a PO-MHP with excitation function $\phi_{ij}(s)=a_{ij}w(s)$ and rank faithfulness. If each latent confounder subprocess, along with all its observed surrogates, satisfies Definition˜3.4, then the causal graph consisting of both observed subprocesses and latent confounders can be identified. In particular, when no latent subprocesses exist, the causal graph is fully identifiable through only Phase I of the algorithm.
@@ 5 Experiments
In our simulation studies, we compare our method against five competitive baselines. SHP (qiao2023structural, ), THP (cai2022thps, ), and NPHC (achab2018uncovering, ) are designed for Hawkes processes and assume no latent subprocesses. Hier. Rank (huang2022latent, ) and RLCD (dong2023versatile, ) leverage rank-based methods for i.i.d. data with linear causal relations. We evaluate our method on two types of synthetic data: event sequences generated by the Hawkes process in Eq. (1), and discrete-time data generated directly from the discrete-time model in Eq. (2). We compare across various synthetic graphs, including the fully observed graph in Fig. 1(b) and five causal structures with latent subprocesses in Figs. 2(a) and 3(a) to 3(d). Performance is measured using F1-score, and the results are the average of ten trials run on a personal PC (CPU). Additionally, we test our method on a larger causal graph and conduct sensitivity analyses on the choice of time interval $\Delta$ and violations of rank faithfulness. Detailed settings and additional results are reported in Appendix˜Q due to page limits. As shown in Fig. 4, our method consistently outperforms baselines across both fully and partially observed graphs. Notably, for graphs with latent subprocesses, a larger sample size is often required, likely because the spectral radius of stationary Hawkes processes is bounded by one, causing causal influences to attenuate through latent subprocesses, thus requiring more data for reliable estimation.
@@
Figure 5: Inferred causal subgraph from the cellular network dataset, where Alarm 8 is successfully identified as a latent subprocess.
@@
We evaluate our method on a real-world dataset—the cellular network dataset (qiao2023structural, ), which contains ground truth relations validated by domain experts. The dataset comprises 18 distinct alarm types and approximately 35,000 alarm events collected over eight months from an operational telecommunication network. For evaluation, we focus on a subgraph involving five alarm types (Alarms 1–4 and 8, with Alarm id=0-3, 7), where Alarm 8 is treated as a latent subprocess by manual exclusion. Notably, both Alarm 2 and Alarm 4 are observed effects of the latent subprocess (Alarm 8), allowing us to infer its presence. The inferred causal graph, shown in Fig. 5, closely matches the ground truth, demonstrating the effectiveness of our approach. Additional details are provided in Section˜Q.4.
@@ 6 Conclusion and Future Work
In this paper, we proposed a principled framework for structure learning in partially observed multivariate Hawkes processes (PO-MHP). By leveraging sub-covariance rank conditions and a carefully designed path constraint, our method effectively identifies both causal relationships among observed subprocesses and latent confounders influencing them. Specifically, we established necessary and sufficient conditions for inferring latent subprocesses and identifying causal relations, and developed a two-phase iterative algorithm with identifiability guarantees to recover the full causal graph. Notably, our approach naturally extends to discrete time series data, given its foundation in the discretized representation of Hawkes processes. Future work includes relaxing the identification conditions to broaden applicability, and applying our method to diverse real-world datasets for deeper domain insights.
@@ Appendix A Related Work
This work is closely related to three areas: point processes, Hawkes processes, and causal discovery methods.
@@
Extensive efforts have been devoted to understanding temporal dependencies in point processes. Meek [36] introduced a graphical framework for general point processes, leveraging $\delta^{*}$-separation and process independence to connect graphical representations with statistical properties. Gunawardana and Meek [19] proposed a one-dimensional point process model with piecewise-constant conditional intensity, utilizing a closed-form Bayesian approach to infer temporal dependencies between event types. Chwialkowski and Gretton [9] developed a kernel-based independence test applicable to general random processes, providing a nonparametric perspective on dependency learning.
Several studies have focused on specific structures within point processes. Basu et al. [4] investigated Granger causality for discrete transition processes while incorporating inherent grouping structures. Daneshmand et al. [13] proposed a continuous-time diffusion network inference method based on a parametric cascade generative process, advancing the modeling of temporal influence. In the context of marked point processes, Didelez [14] introduced a class of graphical models capable of capturing local independence over different marks, offering a more generalized approach to analyzing dependencies in complex systems.
@@
Hawkes processes [20, 33] constitute a class of point processes characterized by self-exciting intensity functions that model how past events affect the likelihood of future events. Much of the research on learning temporal dependencies in Hawkes processes builds upon the concept of Granger causality [18]. Various extensions of the Hawkes process have been developed using predefined excitation functions, including exponential functions [17, 60, 57, 7], power-law functions [59], and nonparametric functions [34, 35].
Regularization techniques have played a critical role in learning optimal temporal dependencies. Xu et al. [55] represented excitation functions using a series of basis functions, applying a sparse-group-lasso regularizer for optimization. Zhou et al. [60] proposed a convex optimization framework incorporating nuclear norm and $L_{1}$-norm regularizations to promote sparsity. Similarly, Ide et al. [25] introduced a cardinality-regularized Hawkes process framework, employing the $L_{0}$-norm to enforce sparsity more rigorously. Nonparametric methods have also been explored; for example, Achab et al. [1] estimated the matrix of integrated kernels in Hawkes processes, while Zhang et al. [58] developed a deep neural network combined with attribution techniques to infer Granger causality. However, these methods primarily address causal relationships among observed processes, overlooking the influence of latent processes.
Kirchner [30] showed that as the time interval decreases, integer-valued autoregressive time series converge to Hawkes processes. Building on this result, Kirchner [29] further approximated Hawkes processes by discretizing them into autoregressive models of bin counts, providing a foundation for analyzing latent processes. This connection is closely related to our work, where we leverage similar approximations to address challenges posed by partial observability and latent influences in Hawkes processes.
Causal Discovery Methods
Causal discovery [37] seeks to uncover underlying causal relationships in observed data and plays a critical role across numerous domains. Traditional methods predominantly focus on data with i.i.d. noise, where the causal structure is assumed to be represented by a directed acyclic graph (DAG). Popular approaches include constraint-based methods such as the PC algorithm [47], score-based methods like GES [8], and function-based methods such as LiNGAM [43].
Latent variables present significant challenges to these methods. To address this, extensions such as the FCI algorithm [48, 46] and its variants [12, 10] leverage conditional independence constraints to infer partial causal structures in the presence of independent latent confounders. Recent advances have extended these methods to handle causally related latent confounders. For example, Huang et al. [22] and Dong et al. [15] proposed methods based on the rank deficiency of the cross-covariance matrix to identify equivalence classes of causal graphs under linear models. Furthermore, Xie et al. [53, 54] and Jin et al. [26] utilized higher-order statistics to accurately identify causal graphs even in the presence of latent confounders.
Several efforts have extended these ideas to time-series data under the i.i.d. noise assumption. For instance, Hyvärinen et al. [24] adapted linear causal models with non-Gaussian noise to time-series settings, while Runge [40] extended the PC algorithm to infer causal structures among lagged variables. Despite these advancements, these methods are inherently designed for data with i.i.d. noise and are not directly applicable to event sequence data with continuous-time dependencies.

### v2

@@ 4.2.2 Characterizing Latent Confounders via Rank Constraints
Consider a PO-MHP with excitation function $\phi_{ij}(s)=a_{ij}w(s)$ under rank faithfulness. The system $\mathcal{N}_{\mathcal{G}}\coloneq\mathcal{O}_{\mathcal{G}}\cup\mathcal{L}_{\mathcal{G}}$ consists of observed subprocesses $\mathcal{O}_{\mathcal{G}}\coloneq\{O_{i}\}_{i=1}^{p}$, and inferred latent confounder processes $\mathcal{L}_{\mathcal{G}}$ whose parent-cause sets remain to be identified. Let $\mathbf{O}_{v}\coloneq\{O_{i}^{(j)}\}^{j\in\{n-m,\dots,n\}}_{i\in\{1,\dots,p\}}$ denote the corresponding observed variable set. For a subprocess $N_{1}\in\mathcal{N}_{\mathcal{G}}$ and a candidate parent-cause set $\mathcal{P}^{\prime}_{\mathcal{G}}\subseteq\mathcal{N}_{\mathcal{G}}$, when either $N_{1}$ is latent, or $\mathcal{P}^{\prime}_{\mathcal{G}}$ contains latent subprocesses, or both, the following condition holds: $\mathcal{P}^{\prime}_{\mathcal{G}}$ is the minimal set such that $\mathrm{rank}\left(\Sigma_{\mathbf{A}_{v},\mathbf{B}_{v}}\right)=|\mathbf{A}_{v}|-1$, where $\mathbf{A}_{v}\coloneq\{\mathcal{D}e(N_{1})^{(j)},\mathcal{D}e(L_{i})^{(j)}\}^{j\in\{n-m,\dots,n\}}_{L_{i}\in\mathcal{P}^{\prime}_{\mathcal{G}}}\cup\{O_{i}^{(j)}\}^{j\in\{n-m,\dots,n-1\}}_{O_{i}\in\mathcal{P}^{\prime}_{\mathcal{G}}}\cup\{O_{i}^{(j)}\}^{j\in\{n-m,\dots,n\}}_{O_{i}\in\mathcal{S}ib(\mathcal{D}e(N_{1}))\,\cup\,\{\mathcal{S}ib(\mathcal{D}e(L_{i}))\}_{L_{i}\in\mathcal{P}^{\prime}_{\mathcal{G}}}}$ and $\mathbf{B}_{v}\coloneq\mathbf{O}_{v}\backslash\left(\mathcal{D}e(N_{1})^{(n)}\cup\{\mathcal{D}e(L_{i})^{(n)}\}_{L_{i}\in\mathcal{P}^{\prime}_{\mathcal{G}}}\right)$ , if and only if $\mathcal{P}^{\prime}_{\mathcal{G}}$ is a subset of the true parent-cause set of $N_{1}$ such that: (i) conditioning on $\mathcal{S}_{\mathcal{G}}\coloneq\mathcal{P}^{\prime}_{\mathcal{G}}\cup\mathcal{D}e(N_{1})\cup\{\mathcal{D}e(L_{i})\}_{L_{i}\in\mathcal{P}^{\prime}_{\mathcal{G}}}\cup\mathcal{S}ib(\mathcal{D}e(N_{1}))\cup\{\mathcal{S}ib(\mathcal{D}e(L_{i}))\}_{L_{i}\in\mathcal{P}^{\prime}_{\mathcal{G}}}$ renders $N_{i}$ locally independent of $\mathcal{N}_{\mathcal{G}}\backslash\mathcal{S}_{\mathcal{G}}$; (ii) for each $L_{i}\in\mathcal{P}^{\prime}_{\mathcal{G}}$, the latent confounder $L_{i}$ with observed effects $\{\mathcal{D}e(N_{1}),\mathcal{D}e(L_{i})\}$ satisfies Definition˜4.4; and (iii) all possible observed surrogates of $N_{i}$ in $\mathcal{O}_{\mathcal{G}}$ have been identified and added into the observed sibling set.
@@
With Theorem˜4.7 (and Proposition˜4.3), we can identify arbitrary causal relations among both observed and inferred latent subprocesses. This naturally raises a final question: How can we further discover new latent subprocesses that are causally related to inferred latent subprocesses, as in Fig. 3(d)? As shown below, the observed surrogate can still be leveraged for this purpose.
@@
Theorem˜4.7 and Theorem˜4.8 are extensions of Proposition˜4.3 and Proposition˜4.5, respectively. These extend the framework by replacing latent subprocesses with their observed surrogates when evaluating the rank of the relevant cross-covariance matrices. Equipped with these four key theorems, we are now ready to present the discovery algorithm in the next section.
@@ 5 Rank-based Discovery Algorithm
In this section, we present a two-phase iterative algorithm that leverages the identification theorems to progressively recover causal relationships among discovered subprocesses and to uncover new latent subprocesses. Let $\mathcal{A}_{\mathcal{G}}$ denote the active process set, consisting of subprocesses whose parent causes remain unidentified. Initially, $\mathcal{A}_{\mathcal{G}}$ is set to the observed subprocess set $\mathcal{O}_{\mathcal{G}}$ and is updated throughout the procedure. Moreover, due to cycles in the summary causal graph, observed subprocesses previously identified as effects may still act as causes for other subprocesses in $\mathcal{A}_{\mathcal{G}}$, and thus remain under investigation. The overall procedure is in Algorithm˜1.
@@
1:  Initialize partial causal graph $\mathcal{G}\coloneq\emptyset$ and active process set $\mathcal{A}_{\mathcal{G}}\coloneq\mathcal{O}_{\mathcal{G}}$.
2:  repeat
3:   $(\mathcal{G},\mathcal{A}_{\mathcal{G}})\leftarrow$ Identifying Causal Relations $(\mathcal{G},\mathcal{A}_{\mathcal{G}},\mathcal{O}_{\mathcal{G}})$. // Phase I
4:   $(\mathcal{G},\mathcal{A}_{\mathcal{G}})\leftarrow$ Discovering New Latent Subprocesses $(\mathcal{G},\mathcal{A}_{\mathcal{G}},\mathcal{O}_{\mathcal{G}})$. // Phase II
5:  until $\mathcal{A}_{\mathcal{G}}$ is empty or no updates occur.
6:  return $\mathcal{G}$
@@
Each iteration begins with Phase I, which identifies the causal structure of under-investigated subprocesses (both observed and latent) in $\mathcal{A}_{\mathcal{G}}$. In this phase, we systematically iterate over each subprocess in $\mathcal{A}_{\mathcal{G}}$ and test its parent causes using the current set $\mathcal{A}_{\mathcal{G}}\cup\mathcal{O}_{\mathcal{G}}$. If a subprocess’s parent-cause set is fully contained by this set, it can be identified using Propositions˜4.3 and 4.7. Once identified, the subprocess is removed from $\mathcal{A}_{\mathcal{G}}$. The phase continues until no further updates occur. Detailed steps are given in Algorithm˜2 in Section˜O.1.
@@
When no additional subprocesses in $\mathcal{A}_{\mathcal{G}}$ can be resolved in Phase I, the algorithm proceeds to Phase II. Here, we search for new latent confounder subprocesses by exhaustively checking all pairs in $\mathcal{A}_{\mathcal{G}}$ using Propositions˜4.5 and 4.8. Identified latent confounders are merged if pairs overlap in subprocesses, implying a shared latent parent cause. $\mathcal{A}_{\mathcal{G}}$ is then updated by adding the new latent subprocesses and removing their effects. After that, the algorithm returns to Phase I. This procedure continues until $\mathcal{A}_{\mathcal{G}}$ is empty or unchanged. Details are provided in Algorithm˜3 in Section˜O.2.
@@
Consider a PO-MHP with excitation function $\phi_{ij}(s)=a_{ij}w(s)$ under rank faithfulness. If every latent confounder subprocess, along with all its observed surrogates $(\geq 2)$, satisfies Definition˜4.4, then the causal graph over observed and latent confounder subprocesses can be identified. When no latent subprocesses exist, the causal graph is fully identifiable using only Phase I.
@@ 6 Experiments
We compare our method against six strong baselines. SHP (Qiao et al., 2023) and THP (Cai et al., 2022) are likelihood-based approaches for discretized (binned) Hawkes data, while NPHC (Achab et al., 2018) is a cumulant-based method for original Hawkes data. Since existing Hawkes-based methods cannot identify latent subprocesses without prior knowledge, we additionally include two rank-based approaches developed for i.i.d. linear latent models, Hier. Rank (Huang et al., 2022) and RLCD (Dong et al., 2023), as well as LPCMCI (Gerhardus and Runge, 2020), a time-series baseline that accounts for exogenous latent confounder variables, though all three rely on strong assumptions not satisfied in our setting (See Appendix˜A). For these methods, we apply them to discretized Hawkes data. For our method, we evaluate both event sequences generated by the Hawkes process in Eq. (1) and data generated directly from the discrete-time model in Eq. (2). We test on six synthetic graph families: the fully observed graph in Fig. 1(b) and five structures with latent subprocesses in Figs. 2(a) and 3(a)3(d). Results are reported as average F1-scores over ten runs on a personal CPU machine. Additional experimental details and extended results (larger graphs, sensitivity to $\Delta$, and robustness to rank-faithfulness violations) are provided in Appendix˜Q. As shown in Fig. 4, our method consistently outperforms all baselines. Notably, latent cases typically require larger sample sizes: because the spectral radius of a stationary Hawkes process is <1<1, causal influences attenuate along latent paths, making reliable detection more data-demanding.
@@
Figure 5: Inferred causal subgraph from the cellular network dataset, where Alarm_id=7 is successfully identified as a latent subprocess.
@@
We evaluate our method on a public cellular network dataset (Qiao et al., 2023) with expert-validated ground truth. The original dataset contains 18 alarm types from 55 devices ($\approx 35\text{k}$ events over eight months), though not every device exhibits all alarms. We focus on device_id=8, which includes the alarms relevant to the subgraph of interest. For evaluation, we consider a five-alarm subgraph (Alarm_ids=0-3 and 7) and treat Alarm_id=7 as latent by manual exclusion. Notably, Alarm_id=1 and Alarm_id=3 are observed effects of this latent subprocess, enabling its recovery from observed data. Our inferred graph (Fig. 5) successfully recovers the latent subprocess and its primary influences; the only discrepancy with the ground truth is a single missing edge, Alarm_id=1 $\to$ Alarm_id=3. Furthermore, on this sub-dataset our method quantitatively outperforms representative baselines; see Section˜Q.4 for details.
@@ 7 Conclusion and Future Work
We presented a principled framework for structure learning in partially observed multivariate Hawkes processes (PO-MHP), capable of uncovering both causal relationships among observed subprocesses and latent confounder subprocesses influencing them. By leveraging sub-covariance rank conditions and a tailored path constraint, we derived necessary and sufficient conditions for identifiability and designed a two-phase iterative algorithm that reconstructs the full causal graph with guarantees. Future work includes relaxing these conditions to broaden applicability and applying to diverse real-world datasets for deeper domain insights.
Ethics Statement
This work makes a primarily methodological contribution, proposing a framework for structure learning in partially observed multivariate Hawkes processes. The study does not involve human subjects, personal or sensitive data, or applications with direct societal or ethical risks. Our experiments are conducted on synthetic data and a publicly available dataset with expert-validated ground truth. We therefore believe that the research does not raise ethical concerns beyond standard scientific integrity, and is fully consistent with the ICLR Code of Ethics.
Reproducibility Statement
We have made significant efforts to ensure the reproducibility of our results. All theoretical results are stated with explicit assumptions, and full proofs are provided in the appendix. Experimental details, including data preprocessing steps, and additional results (e.g., sensitivity to $\Delta$ and robustness analyses), are included in the appendix and supplementary materials. Synthetic data generation procedures are fully specified, and the real-world dataset we use is publicly available.
LLM Usage Statement
Large Language Models (LLMs) were used only as general-purpose tools to aid in polishing the writing and improving clarity of presentation. They were not involved in research ideation, methodological design, experiments, and analysis. No LLM qualifies for authorship.
@@ Appendix A Related Work
We review related work in three areas: point processes, Hawkes processes, and causal discovery.
@@
A large body of research studies temporal dependencies in general point processes. Meek [2014] introduced a graphical framework based on $\delta^{*}$-separation and process independence to connect graphical representations with statistical properties. Gunawardana et al. [2011] proposed a one-dimensional point process model with piecewise-constant conditional intensity, enabling closed-form Bayesian inference of temporal dependencies. Chwialkowski and Gretton [2014] developed a kernel-based independence test for general random processes, offering a nonparametric perspective.
Other works focus on specific structural settings. Basu et al. [2015] studied Granger causality for discrete transition processes with grouping structures. Daneshmand et al. [2014] proposed a parametric cascade generative process for continuous-time diffusion networks. In the setting of marked point processes, Didelez [2008] introduced graphical models capable of capturing local independence across marks, generalizing dependency analysis in complex systems.
@@
Hawkes processes [Hawkes, 1971, Laub et al., 2015] form a prominent class of self-exciting point processes that model how past events influence future occurrences. A major research line estimates Hawkes structures from continuous-time event data, typically via likelihood-based methods. These approaches assume parametric excitation kernels such as exponential or power-law [Zhou et al., 2013, Zhao et al., 2015], or adopt nonparametric procedures [Lewis and Mohler, 2011, Luo et al., 2015], often combined with sparsity or low-rank regularization [Xu et al., 2016, Idé et al., 2021]. The NPHC method [Achab et al., 2018] provides a nonparametric alternative, estimating integrated kernels through moment-matching. While effective when all subprocesses are observed, these approaches do not address the presence of latent components.
When only discretized or binned counts are available, another line of work fits Hawkes models directly from bin data. Shlomovich et al. [2022] proposed an EM algorithm with importance sampling to recover parameters from binned observations without precise timestamps. Qiao et al. [2023] introduced SHP, which learns causal structure from binned event sequences using sparsity-regularized likelihood. Cai et al. [2022] developed THPs, which impose topological constraints to recover causal influences from discrete sequences. These binned-likelihood approaches, however, assume full observability and do not infer the existence or number of latent subprocesses.
Causal Discovery
Causal discovery aims to infer causal relations from data, traditionally under i.i.d. assumptions with DAG structures [Pearl, 2009]. Classical approaches include constraint-based methods (e.g., PC Spirtes et al. [2001]), score-based methods (e.g., GES Chickering [2002]), and functional approaches (e.g., LiNGAM Shimizu et al. [2006]). Latent variables pose major challenges, leading to extensions such as FCI and its variants [Spirtes et al., 1995, Colombo et al., 2012, Claassen et al., 2013], which use conditional independence to infer partial structures under exogenous latent confounder variables.
More recent advances explore settings with causally related latent confounder variables. Huang et al. [2022] and Dong et al. [2023] identify equivalence classes in linear models using second-order (rank) statistics. However, these methods typically rely on restrictive assumptions—such as hierarchical latent structures or cardinality constraints—that are incompatible with Hawkes dynamics. In Hawkes-induced time series, autoregressive representations are dense across lags, observed surrogates are often fewer than effective latent “parents,” and endogenous latent confounder subprocesses arise naturally. Other works, such as Xie et al. [2020, 2022] and Jin et al. [2023], leverage higher-order statistics to improve identifiability, but they still assume i.i.d. data, which may introduce spurious dependencies once temporal dynamics are ignored.
Causal discovery has also been extended to time-series domains. Approaches such as SVAR-based LiNGAM [Hyvärinen et al., 2010] and PC-style temporal methods like PCMCI [Runge, 2020] and LPCMCI [Gerhardus and Runge, 2020] rely on conditional independence tests over lagged variables. Yet these methods presuppose weak autocorrelation and exogenous latent variables, assumptions that are violated in Hawkes processes, which exhibit dense cross-lag interactions and endogenous latent subprocesses.
A.0.1 Detailed Relation to Binned Hawkes Estimation
Shlomovich et al. [2022] address parameter estimation for binned Hawkes processes via a modified EM algorithm when only bin counts $N_{t}=N((t+1)\Delta)-N(t\Delta)$ are observed and exact event times are unavailable. The bin counts are treated as observed data and the unobserved event times $\mathcal{T}$ as latent variables (their Eq. 6). Because direct Monte Carlo sampling of $\mathcal{T}$ is intractable in Hawkes models, they employ importance sampling to simulate within-bin timestamps that match the observed counts, thereby maximizing the (binned) likelihood (see their Sec. 2).
Our goal and methodology differ. Leveraging the link between INAR and linear autoregressive models, Theorem˜4.1 establishes an explicit linear structural representation for discretized multivariate Hawkes processes. This connection enables causal discovery directly over binned variables—including the identification of latent confounder subprocesses—with identifiability guarantees (Propositions˜4.3 and 4.5; Theorems˜4.7 and 4.8). In contrast to likelihood maximization based on simulated event times, our framework uses time-aware rank constraints on cross-covariances to recover causal structure. To the best of our knowledge, prior work has not provided a direct, theoretically grounded reduction from Hawkes processes to linear structural models for the purpose of causal discovery.
A.0.2 Detailed Relation to Rank-Based Latent Discovery
Huang et al. [2022] (and related works by Xie et al. [2022] and Dong et al. [2023]) study latent structure discovery under i.i.d. assumptions and continuous variables. Our problem differs substantively: we aim to recover causal structure among observed and latent subprocesses in multivariate Hawkes processes, where each subprocess is a point process and inference is performed on discretized representations.
Different Data Domain and Causal Assumptions
Huang et al. [2022] (and Xie et al. [2022]) assume a latent hierarchical structure, specifically: (i) there are no direct causal links among observed variables, and all dependencies among observed variables arise exclusively from their latent confounder variables; and (ii) observed variables cannot cause latent variables, i.e., endogenous latent confounders are ruled out (see Eq. 1 and Definition 1 in Huang et al. [2022], and Eq. 1, 2 and Definition 1 in Xie et al. [2022]). Neither assumption is needed in our framework. We allow both direct observed-to-observed edges (see Proposition Proposition˜4.3 in our paper) and the existence of endogenous latent confounder subprocesses that can be caused by observed subprocesses (see Theorem˜4.7 in our paper).
Cardinality Requirements vs. Hawkes Density
Huang et al. [2022], Xie et al. [2022], and Dong et al. [2023] rely on a cardinality condition of the form $|\text{children}|>|\text{parents}|$ for certain latent sets (cf. Definition 4 in Huang et al. [2022], Condition 1 in Xie et al. [2022], Definition 5 in Dong et al. [2023]). This is generally incompatible with discretized Hawkes processes, whose autoregressive representation is inherently dense (Eq. 2 in our paper): if a latent $L_{1}$ causes $O_{2}$, then each discretized variable $O_{2}^{(n)}$ is influenced by many lags of $L_{1}$ (potentially hundreds or thousands in practice), making the required $|\text{children}|>|\text{parents}|$ condition fail systematically. Our method avoids such cardinality assumptions: leveraging the separable excitation (˜1), we place lagged observed variables on both sides of carefully chosen cross-covariance blocks so that rank deficiency reliably signals latent confounders (lines 199–216; Proposition˜4.5; Theorem˜4.8).
Time-Aware vs. i.i.d. Causal Discovery
The above i.i.d. methods do not exploit temporal order and, in principle, can test variables at time $n$ as putative parents of variables at time n1n-1. Our procedure is explicitly time-aware: candidate parents for $t=n$ are restricted to appropriate lags (Propositions˜4.3 and 4.5; Theorems˜4.7 and 4.8), aligning identification with Hawkes dynamics. This distinction mirrors PC [Spirtes et al., 2001] (i.i.d.) vs. PCMCI [Runge, 2020] (time series).